OpenStudy (anonymous):

If (8,b) is a point on the graph of y=5x+3. What is b? b= ( Simplify answer)

OpenStudy (anonymous):

You would plug the b in for x and solve for y i think so it would be y=5(8)+3 so then in would be y= 40 + 3 so y= 43 and y would be the b so b= 43 i think